# UniUIR: Considering Underwater Image Restoration as An All-in-One Learner

### v1

@@ IV-B Qualitative Comparison
As depicted in Tab. II, some early deep learning-based underwater image restoration (UIR) methods such as FUnIE[59] and S-UWnet[60], which employ simpler or shallower networks, yield unsatisfactory results. Interestingly, despite its simplicity, WaterNet [7] performs comparatively well by leveraging preprocessed inputs (white balance, gamma correction) in addition to the original distorted image. More complex network architectures like PUGAN [14], PUIE-Net [61], and Ushape [13] achieve further performance gains. NU2Net [15], which incorporates a URanker model for underwater images into its loss function, also demonstrates strong performance. Thanks to its effectiveness, the proposed UniUIR outperforms recent state-of-the-art (SOTA) UIR methods, including Semi-UIR [26] and HCLR-Net [16]. Moreover, compared to the all-in-one SOTA method PromptIR [31] and the Mamba-based image restoration method MambaIR [43], UniUIR consistently achieves superior results across six metrics on two datasets.
To assess robustness, we also evaluated performance on non-reference datasets (Tab. III). Specifically, all methods are trained on the UIEB-T dataset and then directly tested on the non-reference datasets. In addition to the commonly used non-reference underwater image quality metrics UIQM and UCIQE, we also incorporate the URanker metric. While higher UIQM and UCIQE scores generally suggest better contrast and color vibrancy, these metrics can be biased, potentially overlooking color shifts and artifacts, as highlighted in [15]. The URanker score provides a more comprehensive evaluation of underwater image quality. Although MLLE [25], NU2Net[15], and HCLR-Net [16] tend to achieve higher UCIQE and UIQM scores, Fig. 5 reveals limitations: MLLE introduces chromatic aberrations, while NU2Net and HCLR-Net struggle with accurate color rendition, particularly noticeable in the Challenge-60 [7] and EUVP-330[55] examples. Both Semi-UIR [26] and our proposed UniUIR achieve high UIQM and UCIQE scores, along with strong URanker performance. Importantly, our method explicitly addresses the co-occurrence of multiple distortions in underwater scenes, restoring images in a unified manner. This approach enhances overall perceptual quality while preserving local details, resulting in improved contrast and vibrant colors, ultimately leading to superior performance on these challenging non-reference datasets.
TABLE IV: Effectiveness of components on the T90 [7] dataset. DAM V2 refers to Depth Anything Model V2, MMoEB to Mamba Mixture-of-Experts Block, SFPG to Spatial-Frequency Prior Generator and LCDM to Latent Condition Diffusion Model.
@@ IV-C Quantitative Comparisons
TABLE V: Effectiveness of number of top-k experts on the T90 [7] dataset.
@@
$k$=1 24.71 0.931 0.121 0.597 3.064 1.938
$k$=2 24.94 0.932 0.117 0.602 3.025 1.995
$k$=3 25.11 0.933 0.112 0.601 3.117 2.104
$k$=4 25.01 0.932 0.115 0.607 2.993 2.052
@@ IV-D1 Effects of Different Components
IV-D2 Effects of Number of Top-k Experts
Tab. V presents the effectiveness of varying the number of top-$k$ experts on the T90 dataset. It can be observed that the trends for PSNR, SSIM, LPIPS, and URanker are consistent, all achieving their best values when $k=3$. Although the trends for UCIQE and UIQM differ slightly from the aforementioned metrics, the best UIQM value is also obtained at $k=3$. These findings suggest that selecting $k=3$ experts strikes the optimal balance across all metrics, making it the most suitable configuration for the experiments.
TABLE VI: Performance comparisons on low-light and backlit image enhancement tasks.
@@ IV-D2 Effects of Number of Top-k Experts
IV-E Generalization
IV-E1 Application on Other Low-level Vision Tasks
To validate the effectiveness of the proposed UniUIR for low-light and backlit image enhancement tasks, we retrained UniUIR and compared its performance against state-of-the-art image enhancement methods on the LOL-v1 [63], LOL-v2-real [64], SID [65], and BAID [66] datasets. As shown in Tab. VI, UniUIR achieves state-of-the-art performance across all datasets, outperforming existing methods in terms of PSNR, SSIM, and perceptual quality metrics. This demonstrates the robustness and generalizability of UniUIR in handling complex low-light and backlit scenarios. Fig. 8 provides qualitative results, highlighting our ability to produce enhanced images with corrected illumination and enhanced texture in low-light and backlit scenes. Existing enhancement methods often struggle in scenarios involving intricate lighting conditions or high-frequency textures. UniUIR addresses these challenges through its adaptive feature extraction and spatial-frequency prior generation, ensuring more accurate illumination correction and texture enhancement. These capabilities enable UniUIR to deliver consistently high-quality results, even in challenging regions.
IV-E2 Application on High-level Vision Tasks
@@
We investigate the impact of restored underwater images on downstream tasks, including depth estimation and image segmentation. i) For depth estimation task, we employed the Depth Anything V2 [21] model to extract depth maps from restored images and compared the performance of various UIR methods. As illustrated in Fig. 9 (orange and red boxes in the second row), UniUIR produced depth maps that closely matched those suggested by the reference image, outperforming previous methods in handling depth-related distortions. This indicates that UniUIR better preserves depth consistency and enhances performance on depth estimation tasks. ii) For image segmentation task, we used enhanced images from the T90 dataset produced by various UIR methods and directly applied the Segment Anything Model (SAM) [69] to evaluate the impact of underwater image restoration on segmentation performance. Notably, as shown in Fig. 9 (fourth row, yellow dashed box), only HCLR-Net [16] and our UniUIR correctly identified the background diver’s fins, despite occlusion by foreground fish. However, the visual quality of HCLR-Net’s restoration is inferior to that of UniUIR. Interestingly, in the red dashed box (fourth row, Fig. 9), the reference image’s segmentation erroneously merged two distinct fish into one, while all UIR methods, including the distorted input, successfully distinguished them. This highlights the importance of balancing human-perceived quality and machine vision performance for downstream tasks.
@@
These results demonstrate that UniUIR not only achieves superior visual quality but also significantly improves the utility of restored underwater images for machine vision tasks. The findings underscore the necessity of designing UIR methods that optimize both human-perceived quality and downstream task performance.
@@ IV-F Limitation and Future Work
Although our proposed method consistently outperforms state-of-the-art approaches in underwater image enhancement, certain limitations persist. As illustrated in Fig. 10, existing methods struggle with enhancing underwater images under extreme low-light and blurry conditions. This difficulty can be attributed to the fact that most current datasets emphasize common distortions such as color shifts and low contrast, with limited examples of extreme scenarios like those encountered in low-light environments. Consequently, the learning capacity of models is constrained, leading to challenges in handling more complex conditions. Moreover, our method leverages the pre-trained Depth Anything V2 model, which has been primarily trained on natural scenes containing minimal underwater data. Directly applying its pre-trained weights may result in suboptimal capture of depth features specific to underwater settings.
@@ V Conclusion
In this paper, we systematically review the current challenges in underwater image restoration and propose UniUIR, an all-in-one approach designed to address the complex underwater scenario. To effectively harness the interdependencies among various degradations in underwater environments, we developed the Mamba Mixture-of-Experts module. This module allows each expert to specialize in recognizing specific types of degradation and collaboratively infer task-specific priors while preserving both local and global feature representations. To enhance the model’s ability to identify inconsistencies between distortions across different image regions, we integrate scene depth information using the Depth Anything V2 model. Additionally, we propose a spatial-frequency prior generator that extracts degradation priors from both spatial and frequency domains via task-adaptive routing. This mechanism enables conflicting tasks to utilize distinct network paths, thereby reducing task interference. Compared to state-of-the-art underwater image restoration methods, UniUIR delivers superior results, as evidenced by extensive qualitative and quantitative evaluations.

### v2

@@ IV-B Qualitative Comparison
As shown in Tab. II, early deep UIR methods like FUnIE [56] and S-UWnet [57] with shallow architectures yield suboptimal results. While simpler methods such as WaterNet [4] benefit from preprocessed inputs (e.g., white balance, gamma correction), more advanced architectures (PUGAN [10], PUIE-Net [58], Ushape [9]) and learning strategies (NU2Net [11] with URanker-based loss) achieve better performance. Our UniUIR demonstrates superior performance compared to recent SOTA methods, including Semi-UIR [20], HCLR-Net [12], DiffUIR[60], and AdaIR [61], across six metrics on two datasets.
To assess robustness, we also evaluated performance on non-reference datasets (Tab. III). Specifically, all methods are trained on the UIEB-T dataset and then directly tested on the non-reference datasets. In addition to the commonly used non-reference underwater image quality metrics UIQM and UCIQE, we also incorporate the URanker metric. While higher UIQM and UCIQE scores generally suggest better contrast and color vibrancy, these metrics can be biased, potentially overlooking color shifts and artifacts, as highlighted in [11]. The URanker score provides a more comprehensive evaluation of underwater image quality. Although MLLE [19], NU2Net[11], and HCLR-Net [12] tend to achieve higher UCIQE and UIQM scores, Fig. 5 reveals limitations: MLLE introduces chromatic aberrations, while NU2Net and HCLR-Net struggle with accurate color rendition, particularly noticeable in the Challenge-60 [4] and EUVP-330[50] examples. Both Semi-UIR [20] and our proposed UniUIR achieve high UIQM and UCIQE scores, along with strong URanker performance. Importantly, our method explicitly addresses the co-occurrence of multiple distortions in underwater scenes, restoring images in a unified manner. This approach enhances overall perceptual quality while preserving local details, resulting in improved contrast and vibrant colors, ultimately leading to superior performance on these challenging non-reference datasets. Fig. 7 demonstrates that our method achieves superior UIF and UIQI scores compared to competing approaches across five benchmark datasets (U45, EUVP-330, SQUID-16, Challenge-60, and UCCS). This consistent performance advantage across diverse no-reference evaluation scenarios confirms the robustness and effectiveness of our approach in enhancing underwater imagery without requiring reference images.
TABLE IV: Effectiveness of components on the T90 [4] dataset. DAM V2 denotes Depth Anything Model V2, MMoEB denotes Mamba Mixture-of-Experts Block, SFPG denotes Spatial-Frequency Prior Generator, and LCDM denotes Latent Conditional Diffusion Model.
@@ IV-C Quantitative Comparisons
TABLE V: Effect of number of experts on the T90 [4] dataset.
@@
nn=1 24.71 0.931 0.121 0.597 3.064 1.938
nn=2 24.94 0.932 0.117 0.602 3.025 1.995
nn=3 25.11 0.933 0.112 0.601 3.117 2.104
nn=4 25.01 0.932 0.115 0.607 2.993 2.052
TABLE VI: Effect of the number of MMoEBs on the T90 [4] dataset.
[…]
$[2,4,4,6]$ 23.78 63.41 0.591 3.037 0.476 0.485
$[3,5,6,6]$ 29.19 78.07 0.601 3.117 0.481 0.496
$[4,6,6,8]$ 34.58 91.26 0.604 3.088 0.487 0.504
TABLE VII: Effect of loss function ablation on the T90 [4] dataset.
[…]
w/o $\mathcal{L}_{depth}$ 24.72 0.930 0.592 2.981 0.465 0.477
w/o $\mathcal{L}_{grad}$ 24.93 0.932 0.603 3.069 0.472 0.481
w/o $\mathcal{L}_{diff}$ 24.76 0.929 0.587 3.016 0.467 0.484
All loss 25.11 0.933 0.601 3.117 0.481 0.496
@@ IV-D1 Effects of Different Components
IV-D2 Effects of Number of Experts
Tab. V presents the effectiveness of varying the number of experts on the T90 dataset. It can be observed that the trends for PSNR, SSIM, LPIPS, and URanker are consistent, all achieving their best values when $n=3$. Although the trends for UCIQE and UIQM differ slightly from the aforementioned metrics, the best UIQM value is also obtained at $n=3$. These findings suggest that selecting $n=3$ experts strikes the optimal balance across all metrics, making it the most suitable configuration for the experiments.
IV-D3 Effects of Number of MMoEBs
As shown in Tab. VI, the [3,5,6,6] configuration achieves the best trade-off between performance and efficiency. It yields the highest UIQM score while maintaining moderate model complexity. Although the [4,6,6,8] variant achieves slightly better results in UCIQE, UIF, and UIQI, it incurs significantly higher computational cost. These results suggest that increasing the number of experts does not guarantee consistent performance gains and must be carefully balanced with efficiency constraints.
IV-D4 Effects of Each Loss Term
As shown in Table VII, removing any loss leads to performance drops in most metrics, particularly in PSNR and SSIM, indicating their overall effectiveness. Notably, the model without $\mathcal{L}_{grad}$ achieves a slightly higher UCIQE (0.603 vs. 0.601), suggesting that gradient supervision may slightly over-sharpen structures under this metric. However, the full model achieves the best balance across all metrics, demonstrating that the combined use of $\mathcal{L}_{depth}$, $\mathcal{L}_{grad}$, and $\mathcal{L}_{diff}$ yields the most robust and perceptually favorable restoration.
Figure 9: Expert selection in MMoEM under various degradation types.
TABLE VIII: Comparison of model complexity and time efficiency across different methods in terms of Parameters (M), FLOPs (G), and Inference Time (ms).
[…]
WaterNet [4] 24.81 193.7 314
FUnIE [56] 7.7 10.7 38
Ucolor [8] 157.4 34.68 142
PUIE-Net [58] 10.69 30.09 262
U-shape [9] 65.60 66.20 97
PUGAN [20] 95.66 72.05 276
NU2Net [11] 3.1 10.4 67
Semi-UIR [20] 12.78 36.46 147
PromptIR [24] 32.96 158.14 111
HCLR-Net [12] 4.87 401.97 239
MambaIR [40] 26.78 67.38 133
AdaIR [61] 28.77 147.5 116
UniUIR (Ours) 29.19 78.07 105
TABLE IX: Comparison of performance, model complexity, and efficiency across UniUIR components.
[…]
UniUIRS1 27.95 70.12 84 2.167 0.493 0.511
UniUIRS2-V1 (w/o DAM V2) 29.19 65.56 80 1.996 0.483 0.492
UniUIRS2-V2 (w/o SFPG) 26.41 63.24 93 2.019 0.474 0.487
UniUIRS2-V3 (w/o LCDM) 27.56 68.33 75 2.071 0.472 0.483
UniUIRS2-V4 (w/ MMoE-UIR) 23.37 51.32 61 1.924 0.465 0.478
UniUIRS2 (Ours) 29.19 78.07 105 2.104 0.481 0.496
TABLE X: Comparison of performance and inference time across different settings.
[…]
Baseline 25.11 0.933 0.601 3.117 2.104 0.481 0.496 78.07 105
[…]
TABLE XI: Performance comparisons on low-light and backlit image enhancement tasks.
@@ IV-D4 Effects of Each Loss Term
IV-E Discussion
IV-E1 Performance vs. Efficiency
As summarized in Tab. VIII, UniUIR achieves a favorable trade-off between restoration quality and computational efficiency. It is more efficient than WaterNet, Ucolor, and several GAN-based methods in terms of FLOPs and latency, while outperforming them in visual quality. Although NU2Net runs faster, its significantly lower capacity may limit restoration performance. Compared to recent approaches like PromptIR and AdaIR, UniUIR offers better efficiency and accuracy.
As shown in Tab. IX, the full UniUIR model achieves the best restoration quality at the cost of higher computational load. Notably, UniUIRS1 outperforms UniUIRS2 due to its use of reference image information during inference, which is not available in the latter. The removal of key components, including DAM V2, SFPG, or LCDM, or their substitution with lighter alternatives such as MMoE-UIR, reduces complexity and latency but leads to performance degradation. This result confirms their essential role in maintaining high-quality restoration.
As shown in Tab. X, the results indicate that while increasing the number of diffusion steps can further enhance the restoration quality across various metrics, it also significantly increases the inference time. For instance, the 4-Step→8-Step setting achieves the highest PSNR but requires 154 ms for inference, which is considerably longer than the baseline. Conversely, reducing the number of diffusion steps can drastically reduce the inference time. However, this comes at the cost of slightly degraded performance, with PSNR dropping to 24.75 dB. Interestingly, the DAM V2→Depth-Pro [69] setting offers a balance between performance and efficiency. It not only improves all image quality metrics compared to the baseline but also reduces the inference time. This suggests that integrating more advanced and lightweight depth estimation models into the architecture can effectively improve both restoration quality and computational efficiency.
IV-E2 Expert Selection Analysis
As shown in Fig. 9, $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$ are more readily activated when the primary degradation is color cast, whereas $\mathcal{E}_{1}$ and $\mathcal{E}_{3}$ are more frequently activated under haze degradation. This demonstrates that the model can adaptively and dynamically route different degradation patterns to distinct expert combinations, effectively implementing a decomposition-and-specialization strategy within the network. This adaptive expert activation mechanism ensures that the model does not merely memorize a single restoration strategy but instead utilizes specialized sub-networks to handle specific distortions, thereby enhancing both the interpretability and effectiveness of the restoration process.
IV-E3 Application on Other Low-level Vision Tasks
To evaluate UniUIR on low-light and backlit enhancement, we retrained it and compared with state-of-the-art methods on LOL-v1 [63], LOL-v2-real [64], SID [65], and BAID [66]. As shown in Tab. XI, UniUIR achieves superior performance in PSNR, SSIM, and perceptual quality, demonstrating its robustness and generalizability. Qualitative results in Fig. 10 show that UniUIR effectively corrects illumination and enhances texture, even under complex lighting or high-frequency details. Unlike existing methods, UniUIR leverages adaptive feature extraction and spatial-frequency priors to handle challenging conditions, delivering consistently high-quality enhancement.
IV-E4 Application on High-level Vision Tasks
@@
We evaluate the impact of restored underwater images on downstream tasks: depth estimation and image segmentation. i) For depth estimation, we use Depth Anything V2 [15] to extract depth maps from restored images. As shown in Fig. 11 (second row, orange/red boxes), UniUIR produces results closest to the reference, demonstrating superior depth consistency and distortion correction. ii) For segmentation, we used restored images from the T90 dataset produced by various UIR methods and directly applied the Segment Anything Model (SAM) [62]. As shown in the fourth row (yellow dashed box), both HCLR-Net [12] and UniUIR correctly segment the occluded diver’s fins, while UniUIR yields superior visual quality. Notably, in the red dashed box, the reference image mismerges two fish, while all UIR methods (even the input) correctly separate them, highlighting the need to balance perceptual quality and machine vision performance.
@@
These results show that UniUIR enhances not only visual quality but also the utility of restored images for downstream tasks, advocating for UIR methods that optimize both human and machine perception.
@@ IV-F Limitation and Future Work
Although our proposed method consistently outperforms state-of-the-art approaches in underwater image enhancement, certain limitations persist. As illustrated in Fig. 12, existing methods struggle with enhancing underwater images under extreme low-light and blurry conditions. This difficulty can be attributed to the fact that most current datasets emphasize common distortions such as color shifts and low contrast, with limited examples of extreme scenarios like those encountered in low-light environments. Consequently, the learning capacity of models is constrained, leading to challenges in handling more complex conditions. The scarcity of such samples not only limits the model’s exposure to severe degradation patterns but also biases the training objective toward average-case performance, reducing its robustness in rare yet critical situations. Moreover, our method leverages the pre-trained Depth Anything V2 model, which has been primarily trained on natural scenes containing minimal underwater data. Directly applying its pre-trained weights may result in suboptimal capture of depth features specific to underwater settings.
@@ V Conclusion
In this paper, we systematically review the current challenges in underwater image restoration and propose UniUIR, an all-in-one approach designed to address the complex underwater scenario. To effectively harness the interdependencies among various degradations in underwater environments, we developed the Mamba Mixture-of-Experts module. This module allows each expert to specialize in distinct aspects of degradation and collaboratively infer task-specific priors while preserving both local and global feature representations. To enhance the model’s ability to identify inconsistencies between distortions across different image regions, we integrate scene depth information using the Depth Anything V2 model. Additionally, we propose a spatial-frequency prior generator that extracts degradation priors from both spatial and frequency domains via task-adaptive routing. This mechanism enables conflicting tasks to utilize distinct network paths, thereby reducing task interference. Compared to state-of-the-art underwater image restoration methods, UniUIR delivers superior results, as evidenced by extensive qualitative and quantitative evaluations.